OpenStudy (anonymous):

Find the indicated derivative. In this case, the independent variable is a (unspecified) function of t.

OpenStudy (anonymous):

y= x^1/2 + (1/x^1/2) , x=4 when t=1 , dx/dt t=1 = -8

OpenStudy (anonymous):

your notation looks a bit confusing

OpenStudy (anonymous):

sorry, let me try fixing it. \[y=\sqrt{x} + \frac{ 1 }{ \sqrt{x}} , x=4 when t=1, \frac{ dx }{ dt }| t=1| = -8\]

OpenStudy (anonymous):

sorry, i don't understand this '-8' jazz at the end

OpenStudy (anonymous):

It's okay, I don't either. I have no idea what any of this means!

OpenStudy (anonymous):

anyone?

hartnn (hartnn):

i suppose you need to find dy/dt ?

hartnn (hartnn):

find dy/dx, then use \(\large \dfrac{dy}{dx}=\dfrac{dy}{dt}\dfrac{dt}{dx}\) where, \(\dfrac{dt}{dx}= 1/(\dfrac{dx}{dt})=1/(-8)\)

OpenStudy (anonymous):

I'm not quite sure how I'm supposed to use that

hartnn (hartnn):

could you first find dy/dx ?

OpenStudy (anonymous):

would it be : \[\frac{ .5 }{ x ^{.5} }- \frac{ .5 }{ x ^{1.5} }\] ?

hartnn (hartnn):

yes, thats correct :) now it would be easier if you first plug in x=4 before proceeding to: \(\large \dfrac{dy}{dx}=\dfrac{dy}{dt}\dfrac{dt}{dx}\)

hartnn (hartnn):

\(\large \frac{ .5 }{ 4 ^{.5} }- \frac{ .5 }{ 4 ^{1.5} }=..?\)

OpenStudy (anonymous):

okay found it, .1875? Not sure how to proceed after that

hartnn (hartnn):

yes, now at t=1, its given that dx/dt = -8 we found dy/dx = 0.1875 use \(\large \dfrac{dy}{dx}=\dfrac{dy}{dt}\dfrac{dt}{dx} \implies 0.1875=\dfrac{dy}{dt}*(-8)\) find dy/dt from here.

OpenStudy (anonymous):

\[\frac{ -3 }{ 128 } \] or am I completely lost? Cause according to my assignment that was incorrect

OpenStudy (anonymous):

anyone know what I did wrong?